# Towards deep physics-informed Kolmogorov–Arnold networks

### v1

@@ A.1 Proposed Initialization Scheme
A.2 Chebyshev-based KAN Derivative
Following [23], we consider small activations at initialization and adopt the linear-regime approximation, where
@@ A.2 Chebyshev-based KAN Derivative
Appendix B Implementation Details
All neural network architectures utilized in this study are implemented in JAX [72] using the jaxKAN framework [54] and trained at the highest precision settings on an NVIDIA GeForce RTX 4090 GPU. Their performance is assessed in terms of the L2 error of the predicted solution, $\mathbf{u}_{\text{pred}}$, relative to a reference solution, $\mathbf{u}_{\text{ref}}$, i.e.,
@@ Appendix B Implementation Details
where $\mathcal{E}$ and $\lVert\cdot\rVert_{2}$ denote the relative $L^{2}$ error and $L^{2}$ norm, respectively.
Appendix C Function Fitting Benchmarks
@@
Appendix D Studied Partial Differential Equations
In this appendix, we present the PDEs studied throughout this work, including their governing equations, boundary and/or initial conditions, and corresponding reference solutions. The equations are presented with the specific parameter values used for this study and are listed in the same order in which they appear in Section 5 of the main text.
Allen–Cahn Equation
Figure 16: Reference solution of the Allen–Cahn equation on $t\in[0,1]$, $x\in[-1,1]$.
The Allen–Cahn equation on the spatiotemporal domain $t\in[0,1]$, $x\in[-1,1]$ is given by
$\frac{\partial u}{\partial t}\;-\;10^{-4}\frac{\partial^{2}u}{\partial x^{2}}\;=\;5\,(u-u^{3}).$ (83)
It is considered with initial condition
@@ Appendix D Studied Partial Differential Equations
The reference solution shown in Figure 16 corresponds to the data used in [23] and accessed from the paper’s accompanying GitHub repository [73].
Burgers’ Equation
The viscous Burgers’ equation on the spatiotemporal domain $t\in[0,1]$, $x\in[-1,1]$ is given by
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}\;=\;\frac{1}{100\pi}\frac{\partial^{2}u}{\partial x^{2}}.$ (86)
It is considered with initial condition
@@
The reference solution shown in Figure 17 corresponds to the data used in [23] and accessed from the paper’s accompanying GitHub repository [73].
Figure 17: Reference solution of Burgers’ equation on $t\in[0,1]$, $x\in[-1,1]$.
Korteweg–De Vries Equation
The Korteweg–De Vries equation on the spatiotemporal domain $t\in[0,1]$, $x\in[-1,1]$ is given by
$\frac{\partial u}{\partial t}\;+\;u\,\frac{\partial u}{\partial x}\;+\;0.022^{2}\,\frac{\partial^{3}u}{\partial x^{3}}\;=\;0.$ (89)
It is considered with initial condition
@@
The reference solution shown in Figure 18 corresponds to the data used in [23] and accessed from the paper’s accompanying GitHub repository [73].
Figure 18: Reference solution of the Korteweg–De Vries equation on $t\in[0,1]$, $x\in[-1,1]$.
Sine Gordon Equation
The Sine Gordon equation on the spatiotemporal domain $t\in[0,1]$, $x\in[0,1]$ is given by
@@
It is considered with initial condition
@@
and homogeneous Dirichlet boundary conditions
@@
The analytical solution of this equation is known and given by
$u\left(t,x\right)\;=\;\frac{1}{2}\left[\sin\left(\pi\left(x+t\right)\right)+\sin\left(\pi\left(x-t\right)\right)\right],$ (95)
and is depicted in Figure 19.
Figure 19: Reference solution of the Sine Gordon equation on $t\in[0,1]$, $x\in[0,1]$.
Advection Equation
The advection equation on the spatiotemporal domain $t\in[0,1]$, $x\in[0,2\pi]$ is given by
$\frac{\partial u}{\partial t}\;+\;20\,\frac{\partial u}{\partial x}\;=\;0.$ (96)
It is considered with initial condition
@@
and periodic boundary conditions
@@
The analytical solution of this equation is known and given by
$u\left(t,x\right)\;=\;\sin\left(\text{mod}\left(x-20t,2\pi\right)\right),$ (99)
and is depicted in Figure 20.
Figure 20: Reference solution of the advection equation on $t\in[0,1]$, $x\in[0,2\pi]$.
Helmholtz Equation
@@
$\frac{\partial^{2}u}{\partial x^{2}}\;+\;\frac{\partial^{2}u}{\partial y^{2}}\;+\;u\;=\;\left[1-\pi^{2}\left(a_{1}^{2}+a_{2}^{2}\right)\right]\sin\!\left(a_{1}\pi x\right)\sin\!\left(a_{2}\pi y\right).$ (100)
It is considered with homogeneous Dirichlet boundary conditions
@@
The analytical solution of this equation is known and given by
@@
In Figure 21 we depict this solution for $a_{1}=1$ and $a_{2}=4$.
Figure 21: Analytical solution of the Helmholtz equation on $1,1]^{2}$ with $a_{1}=1$, $a_{2}=4$.
Poisson Equation
@@
$\frac{\partial^{2}u}{\partial x^{2}}\;+\;\frac{\partial^{2}u}{\partial y^{2}}\;=\;-2\pi^{2}\omega^{2}\sin\!\left(\omega\pi x\right)\sin\!\left(\omega\pi y\right).$ (103)
It is considered with homogeneous Dirichlet boundary conditions
@@
The analytical solution of this equation is known and given by
$u(x,y)\;=\;\sin\!\left(\omega\pi x\right)\sin\!\left(\omega\pi y\right).$ (105)
In Figure 22 we depict this solution for $\omega\in\left\{1,2,4\right\}$.
Figure 22: Analytical solution of the Poisson equation on $1,1]^{2}$ with $\omega=1$ (left), $\omega=2$ (middle) and $\omega=4$ (right).
Appendix E PirateNet Architecture

### v2

@@ B.1 Proposed Initialization Scheme
B.2 Chebyshev-based KAN Derivative
Following [PirateNets], we consider small activations at initialization and adopt the linear-regime approximation, where
@@ B.2 Chebyshev-based KAN Derivative
Appendix C Implementation Details
All neural network architectures utilized in this study are implemented in JAX [jax] using the jaxKAN framework [jaxKAN] and trained at the highest precision settings on an NVIDIA GeForce RTX 4090 GPU. Their performance is assessed in terms of the L2 error of the predicted solution, $\mathbf{u}_{\text{pred}}$, relative to a reference solution, $\mathbf{u}_{\text{ref}}$, i.e.,
@@ Appendix C Implementation Details
where $\mathcal{E}$ and $\lVert\cdot\rVert_{2}$ denote the relative $L^{2}$ error and $L^{2}$ norm, respectively. In the following, we provide a breakdown of the training settings and hyperparameter configurations used for each family of experiments presented in the main text.
C.1 Small-Scale Benchmarks
For the function fitting benchmarks of Section 3.2.1, we generate $4\cdot 10^{3}$ input–output samples uniformly over the domain $1,1]^{d}$, with $d$ being the dimensionality of the function. Networks are trained for $2\cdot 10^{3}$ iterations to minimize the $L^{2}$ loss using the Adam [adam] optimizer with a constant learning rate of $10^{-3}$ in full-batch mode. For the final relative $L^{2}$ error evaluation, we use 1,000 uniformly spaced points in [1,1][-1,1] for the one-dimensional task, a $200\times 200$ grid for the two-dimensional tasks, a $30^{3}$ grid for the three-dimensional Hartmann function, and a $10^{5}$ grid for the five-dimensional Sobol g-function.
As far as the small-scale PDE benchmarks of Section 3.2.2 are concerned, $N_{\text{pde}}=2^{12}$ collocation points are used to enforce the differential operator for each PDE. In the Burgers’ case, $N_{\text{bc}}=2^{6}$ points are used for each of the two boundary conditions, together with $N_{\text{ic}}=2^{6}$ points for the initial condition. For Helmholtz, $N_{\text{bc}}=2^{6}$ points are used for each of the four boundary conditions. Collocation points are sampled once from a uniform grid and remain fixed throughout training. Training is performed for $5\cdot 10^{3}$ iterations using the Adam optimizer with a constant learning rate of $10^{-3}$ in full-batch mode.
C.2 Depth-Scaling Experiments
For the depth-scaling experiments of Section 3.3, we utilize all adaptive methods described in A. In particular, adaptive collocation-point resampling is performed with hyperparameters $\delta=1$ and $C=1$, which is a standard choice in the literature [RAD]. The pool of collocation points to enforce the PDE is constructed using $N_{\text{pool}}=400\times 400$ uniformly distributed samples over the spatiotemporal domain $0,1]\times[-1,1]$. Every $2\cdot 10^{3}$ iterations, $N_{\text{pde}}=2^{12}$ points are resampled for training. Alongside this, we apply the RBA method with hyperparameters $\gamma=0.999$ and $\eta=0.01$ as in [RBA], so that resampling follows the probability density function defined in Eq. (57). For the initial condition, a fixed set of $N_{\text{ic}}=2^{6}$ points is used, while RBA weighting is still applied. No collocation points are used to enforce boundary conditions, which are instead incorporated directly into the network architecture [ExactBCs].
In addition, every $10^{3}$ iterations we apply LRA according to Eqs. (51)-(52), with decay parameter $a=0.9$. Causal training is also employed by partitioning the temporal domain into $M=32$ segments and using $\epsilon=1.0$ [PirateNets]. Models are trained for $10^{5}$ iterations using the Adam optimizer with an initial warm-up phase of $10^{3}$ iterations, reaching a learning rate of $10^{-3}$, followed by exponential decay with a factor of $0.9$ every $2\cdot 10^{3}$ iterations.
C.3 Final Experimental Results
For the extensive comparative benchmarks presented in Section 5, the results presented in Sections 5.15.7 correspond to the exact same configuration as detailed in C.2. For the (2+1)-dimensional heat equation of Section 5.8, the pool of collocation points is constructed using $N_{\text{pool}}=50\times 50\times 50$ uniformly distributed samples over the spatiotemporal domain, while all other settings remain unchanged. Finally, regarding the Navier-Stokes equation of Section 5.9, the pool of collocation points consists of 32 uniformly distributed points in the temporal domain and a $64\times 64$ grid in the spatial domain. Furthermore, due to the problem’s sensitivity to the initial condition, we increase the number of collocation points used to enforce the initial condition to $N_{\text{ic}}=2^{12}$. All other training settings match those described previously.
Appendix D Function Fitting Benchmarks
@@
Appendix E Studied Partial Differential Equations
In this appendix, we present the PDEs studied throughout this work, including their governing equations, boundary and/or initial conditions, and corresponding reference solutions. The equations are listed in the same order in which they appear in Section 5 of the main text.
E.1 Allen–Cahn Equation
The (1+1)-dimensional Allen–Cahn equation on the spatiotemporal domain $t\in[0,1]$, $x\in[-1,1]$ is given by
$\frac{\partial u}{\partial t}\;-\;D\frac{\partial^{2}u}{\partial x^{2}}\;=\;5\,(u-u^{3}),$ (84)
where $D$ represents the diffusion coefficient. This parameter controls the width of the interfacial transition layers between the stable phases ($u=\pm 1$); specifically, smaller values of $D$ lead to sharper, stiffer interfaces. In this study, we consider the stiff regime with $D=10^{-4}$. The problem is subject to the initial condition
@@ E.1 Allen–Cahn Equation
The reference solution shown in Figure 9 of the main text corresponds to the data used in [PirateNets] and accessed from the paper’s accompanying GitHub repository [jaxpi].
E.2 Burgers’ Equation
The (1+1)-dimensional viscous Burgers’ equation on the spatiotemporal domain $t\in[0,1]$, $x\in[-1,1]$ is given by
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}\;=\;\nu\frac{\partial^{2}u}{\partial x^{2}},$ (87)
where $\nu$ represents the kinematic viscosity. This parameter governs the balance between nonlinear convection (which tends to steepen gradients into shocks) and diffusion (which smooths them out); lower values of $\nu$ result in sharper shock fronts. In this benchmark, we use $\nu=1/(100\pi)$. It is considered with initial condition
@@ E.2 Burgers’ Equation
The reference solution shown in Figure 10 of the main text corresponds to the data used in [PirateNets] and accessed from the paper’s accompanying GitHub repository [jaxpi].
E.3 Korteweg–De Vries Equation
The (1+1)-dimensional Korteweg–De Vries equation on the spatiotemporal domain $t\in[0,1]$, $x\in[-1,1]$ is given by
$\frac{\partial u}{\partial t}\;+\;u\,\frac{\partial u}{\partial x}\;+\;\lambda^{2}\,\frac{\partial^{3}u}{\partial x^{3}}\;=\;0,$ (90)
where $\lambda$ governs the dispersion of the system. This parameter balances the nonlinear convective steepening ($uu_{x}$) against dispersive spreading ($u_{xxx}$). We consider the low-dispersion regime with $\lambda=0.022$; low values of $\lambda$ allow nonlinearity to dominate initially, causing the wave to steepen into a train of narrow, high-frequency solitons rather than spreading out smoothly. The problem is considered with initial condition
@@ E.3 Korteweg–De Vries Equation
The reference solution shown in Figure 11 of the main text corresponds to the data used in [PirateNets] and accessed from the paper’s accompanying GitHub repository [jaxpi].
E.4 Sine Gordon Equation
The (1+1)-dimensional Sine Gordon equation on the spatiotemporal domain $t\in[0,1]$, $x\in[0,1]$ is given by
@@ E.4 Sine Gordon Equation
This nonlinear hyperbolic PDE is fundamental in relativistic field theory and the modeling of Josephson junctions. We consider the standard dimensionless form, where the characteristic velocity and field mass are normalized to unity. The problem is initialized with
@@
and subject to homogeneous Dirichlet boundary conditions
@@
The analytical solution of this equation, depicted in Figure 12 of the main text, is given by
$u\left(t,x\right)\;=\;\frac{1}{2}\left[\sin\left(\pi\left(x+t\right)\right)+\sin\left(\pi\left(x-t\right)\right)\right].$ (96)
E.5 Advection Equation
The (1+1)-dimensional linear advection equation on the spatiotemporal domain $t\in[0,1]$, $x\in[0,2\pi]$ is given by
$\frac{\partial u}{\partial t}\;+\;c\,\frac{\partial u}{\partial x}\;=\;0,$ (97)
where $c$ represents the advection velocity, determining the speed at which the initial wave profile travels across the domain. We consider a high-velocity regime with $c=20$; such high convection speeds typically pose a challenge for physics-informed learning [failmodes]. The problem is initialized with
@@ E.5 Advection Equation
and subject to periodic boundary conditions
@@
The analytical solution of this equation, depicted in Figure 13 of the main text, is given by
$u\left(t,x\right)\;=\;\sin\left(\text{mod}\left(x-ct,2\pi\right)\right).$ (100)
E.6 Helmholtz Equation
@@
$\frac{\partial^{2}u}{\partial x^{2}}\;+\;\frac{\partial^{2}u}{\partial y^{2}}\;+\;k^{2}u\;=\;\left[k^{2}-\pi^{2}\left(a_{1}^{2}+a_{2}^{2}\right)\right]\sin\!\left(a_{1}\pi x\right)\sin\!\left(a_{2}\pi y\right),$ (101)
where $k=1$ is the wave number and $a_{1},a_{2}$ represent the mode frequencies along the $x$ and $y$ axes, respectively. These parameters control the spatial oscillation density of the solution. The problem is considered with homogeneous Dirichlet boundary conditions
@@
The analytical solution, depicted for $a_{1}=1$ and $a_{2}=4$ in Figure 14 of the main text, is given by
@@
E.7 Poisson Equation
@@
$\frac{\partial^{2}u}{\partial x^{2}}\;+\;\frac{\partial^{2}u}{\partial y^{2}}\;=\;-2\pi^{2}\omega^{2}\sin\!\left(\omega\pi x\right)\sin\!\left(\omega\pi y\right),$ (104)
where $\omega$ governs the spatial frequency of the source term and the solution. Higher values of $\omega$ introduce more rapid oscillations, therefore increasing this parameter allows us to test the network’s ability to resolve high-frequency features. We consider three distinct regimes of increasing complexity: $\omega\in\{1,2,4\}$. The problem is subject to homogeneous Dirichlet boundary conditions
@@
The analytical solution is given by
$u(x,y)\;=\;\sin\!\left(\omega\pi x\right)\sin\!\left(\omega\pi y\right),$ (106)
and is depicted in Figure 15 of the main text.
E.8 Heat Equation
The (2+1)-dimensional multi-scale heat equation on the spatiotemporal domain $t\in[0,1]$, $(x,y)\in[0,1]^{2}$ is given by
$\frac{\partial u}{\partial t}\;-\;D_{x}\frac{\partial^{2}u}{\partial x^{2}}\;-\;D_{y}\frac{\partial^{2}u}{\partial y^{2}}\;=\;0,$ (107)
where $D_{x}$ and $D_{y}$ are the diffusion coefficients along the spatial axes. These parameters govern the rate at which temperature gradients are smoothed out in their respective directions; a larger coefficient leads to rapid dissipation, while a smaller one results in the persistence of high-frequency features. In this benchmark, we introduce extreme anisotropy by setting $D_{x}=(500\pi)^{-2}$ and $D_{y}=\pi^{-2}$. The significant disparity $D_{y}\gg D_{x}$ creates a stiff multi-scale problem where dynamics evolve at radically different rates along the $x$ and $y$ axes. The problem is subject to the initial condition
$u(0,x,y)\;=\;\sin(20\pi x)\sin(\pi y),\quad\text{for }(x,y)\in\Omega,$ (108)
and homogeneous Dirichlet boundary conditions on the boundary of the unit square $\partial\Omega$
$u(t,x,y)\;=\;0,\quad\text{for }(x,y)\in\partial\Omega.$ (109)
The analytical solution, snapshots of which are depicted in Figure 16 of the main text, is given by
$u(t,x,y)\;=\;\sin(20\pi x)\sin(\pi y)\exp\left(-\frac{626}{625}t\right).$ (110)
E.9 Navier-Stokes Equation
The (2+1)-dimensional incompressible Navier-Stokes equations in the velocity-vorticity formulation on the toroidal domain $\Omega=[0,2\pi]^{2}$ for $t\in[0,0.1]$ are given by
$\frac{\partial w}{\partial t}\;+\;\mathbf{u}\cdot\nabla w\;=\;\frac{1}{\text{Re}}\Delta w,\qquad\nabla\cdot\mathbf{u}\;=\;0,$ (111)
where Re is the Reynolds number, $\mathbf{u}=(u,v)$ is the velocity field and $w=\nabla\times\mathbf{u}$ is the vorticity. The Reynolds number characterizes the ratio of inertial forces to viscous forces. Lower values of Re indicate a viscosity-dominated regime where flow features smooth out rapidly, whereas higher values allow for sustained turbulence and complex small-scale structures. We consider a Reynolds number $\text{Re}=100$ and periodic boundary conditions in both spatial directions. Following [Causal, ExpertsGuide], the initial conditions are generated numerically from a random divergence-free velocity field. The reference solution is computed using a high-accuracy pseudo-spectral solver with Crank-Nicolson Runge-Kutta 4 (CN-RK4) time integration on a $64\times 64$ grid (see Data Availability statement). Figure 17 of the main text depicts the resulting reference fields ($u,v$, and $w$) at the final time $t=0.1$.
Appendix F Neural Tangent Kernel Analysis
The Neural Tangent Kernel (NTK) framework serves as an important tool to assess the stability of neural networks by examining the evolution of their gradient descent training dynamics [JacotNTK]. The idea was subsequently extended to PINNs in order to investigate their convergence properties and spectral bias [WangNTK]. Leveraging recent work that derived and analyzed the NTK for Chebyshev-based KANs and cPIKANs [ChebNTK], we utilize this theoretical framework to provide insight into the improved stability offered by our proposed initialization compared to the default scheme.
F.1 Function Fitting
Figure 18: Evolution of the NTK eigenvalue spectra for the small architecture (width 4, depth 3) across the five benchmark functions. The top row displays the spectral evolution under the default initialization, while the bottom row shows the proposed Glorot-like initialization. The solid red line indicates the spectrum at initialization, the green dashed lines represent intermediate checkpoints (every 500 iterations) and the blue dashed line represents the final iteration.
For our NTK-centered analysis, we first use the experimental settings from the function fitting benchmarks in Section 3.2.1. To present the most competitive comparison, we select the training runs where the default initialization achieved its lowest final loss. For these runs, we extract the NTK spectra at initialization and subsequently every 500 iterations until the end of training. The spectra are computed using a randomly selected subset of $2^{8}$ points from the training data. The evolution of the NTK eigenvalues is visualized in Figure 18 for the small architecture and in Figure 19 for the large architecture. In both figures, the top row displays the spectra under the default initialization, while the bottom row corresponds to the proposed Glorot-like scheme.
Figure 19: Evolution of the NTK eigenvalue spectra for the big architecture (width 16, depth 5) across the five benchmark functions. The top row displays the spectral evolution under the default initialization, while the bottom row shows the proposed Glorot-like initialization. The solid red line indicates the spectrum at initialization, the green dashed lines represent intermediate checkpoints (every 500 iterations) and the blue dashed line represents the final iteration.
Across both small and large architectures, the proposed initialization demonstrates superior spectral properties compared to the default scheme. With the exception of $f_{5}$, where the benchmark’s complexity limits the impact of initialization for the studied networks, a consistent pattern emerges: spectra under the default initialization tend to systematically decrease or collapse during training. In contrast, the proposed scheme exhibits the opposite, favorable evolution, expanding to a robust profile. Notably, this stabilization often occurs rapidly; the significant overlap between intermediate and final spectral states (green and blue lines, respectively) indicates that the proposed initialization allows the network to quickly locate a stable optimization path. The distinction in stabilization between the two initialization schemes is particularly evident in the large architecture, where the default initialization yields spectra dominated almost exclusively by very large eigenvalues. Conversely, the NTK for the proposed initialization evolves into a characteristic power-law distribution that spans several orders of magnitude – retaining high eigenvalues for dominant features while also accessing very low values near zero – thereby allowing effective learning across multiple scales.
These observations align closely with the final relative $L^{2}$ error for these runs reported in Table 29. The correspondence is most striking in the large architecture configurations for functions $f_{2}$ and $f_{4}$: under the default initialization, the large networks fail to converge, yielding relative errors exceeding unity. This divergence is a direct consequence of the ill-conditioned spectra observed in the top row of Figure 19, where the dominance of excessively large eigenvalues hinders effective gradient descent. In stark contrast, the proposed initialization not only prevents this divergence but achieves relative errors of $\mathcal{O}(10^{-2})$ or lower, confirming that the induced spectral stability is a prerequisite for accurate function approximation in deeper KAN architectures.
Function Width 4, Depth 3 Width 16, Depth 5
Default Proposed Default Proposed
$f_{1}(x)$ $5.98\times 10^{-2}$ $\mathbf{6.45\times 10^{-3}}$ $2.01\times 10^{-2}$ $\mathbf{1.37\times 10^{-3}}$
$f_{2}(x_{1},x_{2})$ $7.34\times 10^{-2}$ $\mathbf{1.06\times 10^{-2}}$ $1.29\times 10^{0}$ $\mathbf{1.36\times 10^{-2}}$
$f_{3}(x_{1},x_{2})$ $2.19\times 10^{-2}$ $\mathbf{5.17\times 10^{-3}}$ $5.75\times 10^{-1}$ $\mathbf{4.06\times 10^{-3}}$
$f_{4}(x_{1},x_{2},x_{3})$ $9.99\times 10^{-1}$ $\mathbf{5.17\times 10^{-2}}$ $1.37\times 10^{0}$ $\mathbf{4.17\times 10^{-2}}$
$f_{5}(x_{1},\dots,x_{5})$ $9.47\times 10^{-1}$ $\mathbf{9.08\times 10^{-1}}$ $9.45\times 10^{-1}$ $\mathbf{7.50\times 10^{-1}}$
Table 29: Final relative $L^{2}$ errors for the representative training runs used in the NTK analysis for function fitting. The proposed initialization achieves the lowest error across all cases (highlighted in bold).
F.2 Forward PDE Problems
We continue our analysis with the PDE benchmarks of Section 3.2.2. Unlike the function fitting tasks, the loss function here is a composite of multiple terms, therefore we compute separate NTKs for the residuals associated with the PDE operator and the boundary conditions [WangNTK]. As before, we analyze the training runs where the default initialization yielded the lowest final loss and extract the NTK spectra at initialization and subsequently every 1000 iterations until the end of training. The spectra are computed using a subset of $2^{8}$ points for the PDE term and $2^{5}$ points for the boundary condition term. Figure 20 and Figure 21 visualize the spectral evolution for the small (width 4, depth 3) and large (width 16, depth 5) architectures, respectively.
Figure 20: Evolution of the NTK eigenvalue spectra for the small architecture (width 4, depth 3) on PDE benchmarks. The top row displays the spectral evolution under the default initialization, while the bottom row shows the proposed Glorot-like initialization. The solid red line indicates the spectrum at initialization, the green dashed lines represent intermediate checkpoints (every 1000 iterations) and the blue dashed line represents the final iteration.
Figure 21: Evolution of the NTK eigenvalue spectra for the large architecture (width 16, depth 5) on PDE benchmarks. The top row displays the spectral evolution under the default initialization, while the bottom row shows the proposed Glorot-like initialization. The solid red line indicates the spectrum at initialization, the green dashed lines represent intermediate checkpoints (every 1000 iterations) and the blue dashed line represents the final iteration.
The results for the large architecture offer the most critical insight into the stability gap discussed in Section 3.2.2. Under the default initialization (top row), the eigenvalues at initialization are pathologically large. This is most extreme for the Helmholtz PDE term, where the maximum eigenvalue exceeds $10^{20}$. Such excessive magnitudes at initialization force the optimizer to drastically compress the spectrum to fit the target zero residuals, resulting in the observed spectral collapse (the blue line falls far below the red line). This behavior indicates that the network is primarily struggling to suppress its own initial scale rather than learning the solution structure. In contrast, the proposed initialization (bottom row) starts optimization in a much more favorable regime, with maximum eigenvalues typically between $10^{2}$ and $10^{5}$. Similar to the function fitting observations, the spectra under the proposed scheme tend to expand or stabilize rather than collapse. Furthermore, particularly for the Helmholtz equation, the proposed initialization leads to rapid convergence to a stable spectral state, as evidenced by the significant overlap between the intermediate (green dashed) and final (blue dashed) spectra.
Complementing the spectral analysis, Table 30 presents the final relative $L^{2}$ errors for these representative runs. In every case, regardless of architecture size or PDE type, the difference between the two schemes is effectively the difference between divergence and convergence. Models trained under the default initialization systematically fail to learn, yielding relative errors near unity or higher, which aligns perfectly with the spectral collapse observed in the NTK spectra. Conversely, the proposed initialization consistently enables convergence, leading to lower relative $L^{2}$ errors.
PDE Width 4, Depth 3 Width 16, Depth 5
Default Proposed Default Proposed
Burgers $1.00\times 10^{0}$ $\mathbf{4.51\times 10^{-1}}$ $1.05\times 10^{0}$ $\mathbf{3.57\times 10^{-1}}$
Helmholtz $1.03\times 10^{0}$ $\mathbf{4.77\times 10^{-1}}$ $1.14\times 10^{0}$ $\mathbf{1.21\times 10^{-1}}$
Table 30: Final relative $L^{2}$ errors for the representative training runs used in the NTK analysis for forward PDEs. The proposed initialization is the only one achieving convergence across all cases (highlighted in bold).
Appendix G PirateNet Architecture